# Steady state diagram of interacting fermionic atoms coupled to dissipative cavities

### v1

@@ V.1.3 Many-body cooling processes
It is important to note that the cooling effects can only arise due to the fact that in our method the temperature is self-consistently determined from the perturbative effect of the cavity fluctuations (see Sec. III.2). Such phenomena would not be captured by imposing a fixed temperature for the atomic state.
@@ V.1.4 Fluctuation-induced bistability
Figure 16: Simplified phase diagrams varying the atoms-cavity coupling $\hbar g/J$ and (a) the on-site interaction $U/J$ at $\hbar\delta/J\!=\!5$, $\hbar\varGamma/J\!=\!3$; (b) the pump-cavity detuning $\hbar\delta/J$ at U/J=40U/J\!=\!40, $\hbar\varGamma/J\!=\!5$; (c) the dissipation rate $\hbar\varGamma/J$ at $\hbar\delta/J\!=\!5$, U/J=40U/J\!=\!40 in the thermodynamic low hopping limit $L\!\to\!\infty$ at quarter filling. The red line marks the onset of the bistability approximated by Eq. (41). The lower black line in the purple region marks the critical coupling $\hbar g^{\text{MF}}_{\text{cr}}$ obtained for the T=0T\!=\!0 MF approach. Panels (a), (b) are adapted from Ref. Tolle et al. (2025a).
@@
Figure 17: Solution of the system of equations Eqs. (29) with fixed values of $k_{B}T=1/\beta$ varying $\hbar g/J$ at U/J=40U/J\!=\!40, $\hbar\delta/J\!=\!8$, $\hbar\varGamma/J\!=\!1$, $k_{B}T_{\text{max}}/J\!=\!25$ in the thermodynamic limit at quarter filling. (adapted from Ref. Tolle et al. (2025a))
Figure 18: Individual terms of Eq. (29a) and full equation (black line) at $\hbar g/J\!=\!12$, U/J=40U/J\!=\!40, $\hbar\delta/J\!=\!8$, $\hbar\varGamma/J\!=\!1$ in the thermodynamic limit at quarter filling. The system of equations has two stable (blue points, $\lambda_{1(2)}$) and one unstable (red point) solutions at the zero crossings of the black line. Contributions from second line terms in Eq. (29a) with p=1p\!=\!-1 are shown as green lines (close to 0).
Importantly, the bistability does not appear in the T=0T\!=\!0 MF approach (black line, Fig. 15). The crucial ingredient is self-consistent determination of the effective temperature. In particular, the effective temperature in the two solutions differs considerably. We find that the second solution arises with the lower effective temperature. We verified that if the temperature is fixed externally, and not determined by the energy flow between the atoms and the cavity, the bistability is not present. We see in Fig. 17 that regardless of the value of the externally fixed temperature a single stable solution is obtained. Thus, the origin of the fluctuation-induced bistability is fundamentally different from the ones observed in T=0T\!=\!0 MF methods, which are typically caused by steady states of different nature that are coexisting Ritsch et al. (2013), e.g. around a first-order dissipative phase transition Larson et al. (2008b); Minganti et al. (2018); Helson et al. (2022); Ferri et al. (2021), or due to spontaneously breaking of a weak symmetry Baumann et al. (2010); Keeling et al. (2010); Mivehvar (2024). We observe a mean-field bistability for the considered model at half filling (n=1n\!=\!1), which we discuss in Sec. V.2.
@@
The width of the bistability increases approximately linearly with $U/J$ as can be seen in Fig. 16 (a). At fixed $\hbar\delta/J$ the interaction strength controls the position of the resonances between the excited states from which the bistability originates. At large $U/J$ by analyzing Eq. (29) we derive an approximate expression for lower onset of the bistability $g_{\text{bi},1}$ (see Supplemental Material of Ref. Tolle et al. (2025a)). In this limit, the chemical potential scales approximately linearly with the temperature, i.e. $\mu\!\approx\!-\log(2)k_{B}T$. Using this expression for the chemical potential, the self-consistency equation can be simplified and solved for the temperature. Both expressions can be inserted into Eq. (29a). Analyzing this equation further by keeping only the terms which are important around the lower onset of the bistability and using Padé-approximations Baker and Graves-Morris (1996), we find that the onset follows approximately (see Supplemental Material of Ref. Tolle et al. (2025a))
@@ V.2 Half filling (n=1n\!=\!1)
In the absence of the pump the Fermi Hubbard system at half filling can be in a liquid or a Mott-insulating phase depending on the interaction strength. Switching on the pump, a finite density imbalance can only be achieved with an increased density of double occupancies in the classical limit on the low potential sites (see Sec. IV for details). We expect that across the self-ordering transition an initial Mott-insulating state is destroyed and the character of the fermionic state changes to a state with finite density imbalance in the steady state. Taking fluctuations in the coupling perturbatively into account we still find a bistability, i.e. a region of pumping strengths for which two or more steady states with a distinct finite photon number exist. However, the nature of this bistability at half filling is fundamentally different from the bistability we introduced in Sec. V.1.4. Here the nature of the ground state of the effective model, changes its character due to a closing of the gap to the lowest excitations from a Mott-insulator to a charge density wave. Therefore, the bistability is also present at the T=0T\!=\!0 MF level. Such bistabilities are typically associated with a dissipative first-order transition Larson et al. (2008b); Minganti et al. (2018).
@@ V.2.1 Self-ordering transition and bistable solutions
Figure 19: (a) Sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$, (b) inverse temperature $\beta$, (c) scaled photon number and (d) average number of double occupancies as a function of atoms-cavity coupling $\hbar g/J$ at $\hbar\delta/J\!=\!4$; of a finite size system L=6L\!=\!6 at half filling for stable (blue) and unstable (red) results. The T=0T\!=\!0 MF solution is shown as black solid line. The parameters used are U/J=2U/J\!=\!2, $\hbar\varGamma/J\!=\!1$. Vertical lines denote the critical coupling $\hbar g_{\text{cr}}/J$ and resonances in the atomic limit. The orange dashed lines are the approximate scalings for U/J=0U/J\!=\!0 [Eqs. (39)].
Refer to caption(a)
Refer to caption(b)
Figure 20: (a) Sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (b) inverse temperature $\beta$, as a function of atoms-cavity coupling $\hbar g/J$ at $\hbar\delta/J\!=\!4$; of a finite size system L=6L\!=\!6 at half filling for stable (blue) and unstable (red) results. The T=0T\!=\!0 MF solution is shown as black solid line. The parameters used are U/J=8U/J\!=\!8 and $\hbar\varGamma/J\!=\!1$. Vertical lines denote the critical coupling $\hbar g_{\text{cr}}/J$ and resonances in the atomic limit. The orange dashed line is the approximate scaling for U/J=0U/J\!=\!0 [Eqs. (39)].
To start we discuss the results of the T=0T\!=\!0 MF results. We show an example of the self-organization transition for intermediate interaction strength U/J=2U/J\!=\!2 in Fig. 19 and for larger interaction strength U/J=8U/J\!=\!8 in Fig. 20, where we show with the black line the result of the T=0T\!=\!0 MF method. Within the T=0T\!=\!0 MF method, the photon number, the density imbalance and the average value of the double occupancies exhibit a large value in the self-organized solution at the transition for both shown values of $U/J$. This stems from the property that at commensurate filling a finite density imbalance coincides with the creation of double-occupancies on the low-potential sublattice. At larger interaction strength, U/J=8U/J\!=\!8, the emergence of a bistable behavior around the self-ordering transition can be seen in Fig. 20 for T=0T\!=\!0 MF. The stable solution with homogeneous density distribution and zero photon field and the self-ordered steady state above a critical coupling $g^{\text{MF}}_{\text{cr}}$ (black points) coexist and are connected by an unstable branch (gray points). For the plot shown here, the gray line is partly covered by the red curve. The sublattice imbalance and the double occupancy approach their maximal value and the photon number continues to rise very steeply with the pump strength. At intermediate interaction strength U/J=2U/J\!=\!2 a bistability also seems to occur, but the region of the instability seems to be much more narrow (Fig. 19). Due this very narrow bistable region it becomes numerically difficult to identify it unambiguously.
@@
Figure 21: Simplified phase diagrams varying the atoms-cavity coupling $\hbar g/J$ and (a) the on-site interaction $U/J$ at $\hbar\delta/J=4$, $\hbar\varGamma/J\!=\!1$; (b) the pump-cavity detuning $\hbar\delta/J$ at U/J=2U/J\!=\!2, $\hbar\varGamma/J\!=\!1$; (c) the dissipation rate $\hbar\varGamma/J$ at $\hbar\delta/J\!=\!4$, U/J=8U/J\!=\!8 of a finite size system L=6L\!=\!6 at half filling. The lower black line marks the transition between normal [purple region] and the high-$\lambda$ solution in the self-organized phase [blue region] determined by solving Eq. (5) and Eq. (15) with ED. The lower green line marks the critical coupling $\hbar g^{\text{MF}}_{\text{cr}}$ obtained for the T=0T\!=\!0 MF approach. The dashed lines are approximations critical couplings $g_{\text{cr}}^{\text{MF}}$ [Eq. (42)].
@@
The narrowing at small interaction strength of the parameter region where two stable solutions are present can be seen varying the on-site interaction $U/J$ and atoms-cavity coupling $\hbar g/J$ in Fig. 21 (a), marked for the T=0T\!=\!0 MF results by the green hatched region. At small interaction strength it is very hard to identify whether a bistability is present. In contrast, at large interaction strength a broad bistable region occurs. We would like to emphasize that the nature of the bistabilty at half filling differs from the fluctuation-induced bistability at low filling. The onset of the bistability here coincides with the self-organization transition. We attribute this to the change of the nature of the ground state of the effective ionic Hubbard model from the Mott-insulating state to a charge density wave Larson et al. (2008a). This transition we expect around $2\hbar g\lambda\!\approx\!U$ for large $U$ and large $g\lambda$ (as discussed in Sec. IV), where $\lambda$ corresponds to the stable self-ordered solution. The self-consistency condition assuming maximal density imbalance is given by $\langle\hat{\Delta}\rangle\!=\!N$. The resulting critical coupling scales approximately as
@@
(magenta dashed lines in Fig. 21). Here, the dependence on the detuning and the dissipation enters as already seen at lower fillings. The approximate scaling with the $\sqrt{U}$ agrees roughly with the scaling obtained numerically in Fig. 21 (a). For the commensurate filling we cannot make the assumption of a small cavity field. Therefore, the considerations made for the critical coupling strength at lower fillings in Sec. V.1.1 do not apply here.
Another approach to determine the bistability edges involves the stability condition derived in Eq. (25). Assuming the connecting solution connecting low-and high stable solution to be unstable, the coupling for which the condition has a second order zero is roughly at the two critical couplings. We explored this approach numerically and find reasonable agreement with our previously obtained results from solving Eq. (5) and Eq. (15) via ED.
@@
Figure 22: (a) Sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$, (b) inverse temperature $\beta$, (c) scaled photon number and (d) average number of double occupancies as a function of $\hbar\delta/J$ at $\hbar g/J\!=\!4$ of a finite size system L=6L\!=\!6 at half filling for stable (blue) and unstable (red) results. The T=0T\!=\!0 MF solution is shown as black solid line. The parameters used are U/J=2U/J\!=\!2, $\hbar\varGamma/J\!=\!1$. Vertical lines denote the critical coupling $\hbar g_{\text{cr}}/J$ and resonances. The orange dashed lines are the scalings of the analytical approximations for U/J=0U/J\!=\!0 [Eqs. (39)].
Refer to caption(a)
Refer to caption(b)
Figure 23: (a) Sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (b) inverse temperature $\beta$, as a function of $\hbar\varGamma/J$ at $\hbar g/J\!=\!6$ of a finite size system L=6L\!=\!6 at half filling for stable (blue) and unstable (red) results. The T=0T\!=\!0 MF solution is shown as black solid line. The parameters used are U/J=8U/J\!=\!8, $\hbar\delta/J\!=\!4$. The vertical lines denote $2\hbar g\lambda\!=\!U\!+\!\hbar\delta$ (purple dashed) and $\hbar g_{\text{cr}}/J$ (green dashed). The orange dashed line is the scaling of the analytical approximation for U/J=0U/J\!=\!0 [Eqs. (39)].
Refer to caption(a)
Refer to caption(b)
Figure 24: (a) Sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (b) inverse temperature $\beta$ as a function of atoms-cavity coupling $\hbar g/J$ and on-site interaction $U/J$ of a finite size system L=6L\!=\!6 at half filling. The parameters used are $\hbar\delta/J\!=\!4$, $\hbar\varGamma/J\!=\!1$. The hatched region marks the fluctuation-induced bistability region at the transition to the self-organized phase. The lines denote $U\!=\!\hbar\delta$ (red dashed) and $\hbar g_{\text{cr}}/J$ (green dashed).
@@
Figure 25: (a) Sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$, (b) inverse temperature $\beta$, (c) scaled photon number and (d) average number of double occupancies as a function of atoms-cavity coupling $\hbar g/J$ and detuning $\hbar\delta/J$ of a finite size system L=6L\!=\!6 at half filling. The parameters used are U/J=2U/J\!=\!2, $\hbar\varGamma/J\!=\!1$. The hatched region marks the fluctuation-induced bistability region at the transition to the self-organized phase. The lines denote $U\!=\!\hbar\delta$ (red dashed) and $\hbar g_{\text{cr}}/J$ (green dashed).
@@
We turn to to the dependence with the dissipation strength $\hbar\varGamma/J$. For low values of $\hbar\varGamma/J$, the shift of the self-ordering transition between the T=0T\!=\!0 MF method and the approach including fluctuations is relatively small [black and green hatched region in Fig. 21 (c)]. This is due to the fact that in this regime the self-consistently determined temperature is small compared to the excitation gap, which is proportional to $U/J$. Thus, the finite self-consistent temperature does not introduce a significant shift. However, increasing the dissipation $\hbar\varGamma/J$ the width of the bistability region drastically decreases when increasing $\hbar\varGamma/J$ including the fluctuations, while for the T=0T\!=\!0 MF result it even increases in width [as seen in Fig. 21 (c)]. We attribute the shrinking of the bistability region mostly to the rise in the effective temperature when increasing $\varGamma$ [see blue points in Fig. 23 (b)]. In the presence of the fluctuations, from $\hbar\varGamma_{\text{cr}}/J\!\sim\!18.3$ within the numerical accuracy no bistability can be detected anymore, hinting towards a continuous transition. This result resembles the finding in the model considered in Foss-Feig et al. (2017), where, the ordering transition shows the properties of a discontinuous first-order phase transition which turns continuous in the large dissipation limit.
@@
We want to comment on how some of the cooling resonances discussed in detail in Sec. V.1.3 also show up at half filling. The effects are mostly analogous, i.e. varying $\delta$, we note a drastic increase of the photon number around the cooling resonance $U\!=\!\hbar\delta$ [see Fig. 25 and Fig. 22 (c) around the red dashed line]. Here the energy change when creating a double occupancy is resonant to the process of adding a photon to the cavity field.

### v2

@@ V.1.3 Many-body cooling processes
It is important to note that the cooling effects can only arise due to the fact that in our method the temperature is self-consistently determined from the perturbative effect of the cavity fluctuations (see Sec. III.2).
@@ V.1.4 Fluctuation-induced bistability
Figure 16: Simplified phase diagrams varying the atoms-cavity coupling $\hbar g/J$ and (a) the on-site interaction $U/J$ at $\hbar\delta/J\!=\!5$, $\hbar\varGamma/J\!=\!3$; (b) the pump-cavity detuning $\hbar\delta/J$ at U/J=40U/J\!=\!40, $\hbar\varGamma/J\!=\!5$; (c) the dissipation rate $\hbar\varGamma/J$ at $\hbar\delta/J\!=\!5$, U/J=40U/J\!=\!40 in the weak tunneling perturbation in the thermodynamic limit $L\!\to\!\infty$ at quarter filling. The green line marks the transition between normal (purple region) and self-organized phase (blue region) at $\hbar g_{\text{cr}}/J$ determined by solving Eqs. (5) & (15), the solid black line in the purple region marks the critical coupling $\hbar g^{\text{MF}}_{\text{cr}}$ obtained for the T=0T\!=\!0 MF approach solving Eq. (5) via ED. The black lines bordering the hatched region mark the couplings $\hbar g_{\text{bi,1(2)}}$ at finite temperature. The red line marks the onset of the bistability approximated by Eq. (41).
@@
Figure 17: Solution of the system of equations Eqs. (29) with color-coded fixed values of $k_{B}T=1/\beta$ varying $\hbar g/J$ at U/J=40U/J\!=\!40, $\hbar\delta/J\!=\!8$, $\hbar\varGamma/J\!=\!1$, $k_{B}T_{\text{max}}/J\!=\!25$ in the weak tunneling perturbation in the thermodynamic limit at quarter filling. (adapted from Ref. Tolle et al. (2025a))
Figure 18: Individual terms of Eq. (29a) and full equation (black line) at $\hbar g/J\!=\!12$, U/J=40U/J\!=\!40, $\hbar\delta/J\!=\!8$, $\hbar\varGamma/J\!=\!1$ in the weak tunneling perturbation in the thermodynamic limit at quarter filling. The system of equations has two stable (blue points, $\lambda_{1(2)}$) and one unstable (red point) solutions at the zero crossings of the black line. Contributions from second line terms in Eq. (29a) with p=1p\!=\!-1 are shown as green lines (close to 0).
Importantly, the bistability does not appear in the T=0T\!=\!0 MF approach (black line, Fig. 15). The crucial ingredient which causes this effect are the quantum fluctuations in the coupling term between the atoms and the cavity field which act on top of the mean-field decoupling. The fluctuations imprint the dissipative nature of the cavity field onto the atoms, resulting in a heating of the atoms to a finite temperature mixed state. The effective temperature to which the system stabilizes is determined by the complex interplay of the cavity energy scales and the spectrum of the effective atomic Hamiltonian. Whereas at low interaction strength we obtain that the effective temperature is uniquely determined for each coupling strength, in the strongly interacting regime there exists regions where multiple solutions appear. The effective temperature and the strength of the cavity field in these stable solutions differ considerably. We find that the second solution arises with the lower effective temperature. We verified that if the temperature is fixed externally, and not determined by the energy flow between the atoms and the cavity, the bistability is not present. We see in Fig. 17 that regardless of the value of the externally fixed temperature a single stable solution is obtained. Thus, the origin of the fluctuation-induced bistability is fundamentally different from the ones observed in T=0T\!=\!0 MF methods, which are typically caused by steady states of different nature that are coexisting Ritsch et al. (2013), e.g. around a first-order dissipative phase transition Larson et al. (2008b); Minganti et al. (2018); Helson et al. (2022); Ferri et al. (2021), or due to spontaneously breaking of a weak symmetry Baumann et al. (2010); Keeling et al. (2010); Mivehvar (2024). We observe a mean-field bistability for the considered model at half filling (n=1n\!=\!1), which we discuss in Sec. V.2.
@@
The width of the bistability increases approximately linearly with $U/J$ as can be seen in Fig. 16 (a). At fixed $\hbar\delta/J$ the interaction strength controls the position of the resonances between the excited states from which the bistability originates.
At large $U/J$ by analyzing Eq. (29) we derive an approximate expression for lower onset of the bistability $g_{\text{bi},1}$ (see Supplemental Material of Ref. Tolle et al. (2025a)). In the limit $\beta U\!\gg\!1$, we find from Eq. (29b) that fixing the particle number the chemical potential scales approximately linearly with the temperature, i.e. $\mu\!\approx\!-\log(2)k_{B}T$. Using this expression for the chemical potential, the self-consistency equation Eq. (29c) can be simplified and solved for the temperature.
Both expressions can be inserted into Eq. (29a). Analyzing this equation further by keeping only the terms which are important around the lower onset of the bistability and using Padé-approximations around the point $\lambda\!=\!\delta g\tanh(2)/(\varGamma^{2}/4\!+\!\delta^{2})$ Baker and Graves-Morris (1996) we find that the onset follows approximately (see Supplemental Material of Ref. Tolle et al. (2025a))
@@
A direct comparison of the analytic approximation (red line in Fig. 16) and the numerical results (lower black line in Fig. 16) shows a very good agreement of both results. The increase of $g_{\text{bi},1}$ with the interaction strength is well reproduced in Fig. 16 (a) and only small deviations at lower value of $U/J$ are found. Also the increase of $g_{\text{bi},1}$ with the detuning in Fig. 16 (b) is well reproduced by the approximation, and only deviations occur for large $\delta$. Additionally, the dependence on the dissipation strength is very well covered by the approximation and almost no deviations are seen.
@@ V.2 Half filling (n=1n\!=\!1)
In the absence of the pump beam realizing the coupling to the cavity, the Fermi Hubbard system at half filling can be in a liquid or a Mott-insulating phase depending on the interaction strength. Switching on the coupling, a finite density imbalance can only be achieved with an increased density of double occupancies in the semiclassical limit on the low potential sites (see Sec. IV for details). We expect that across the self-ordering transition an initial Mott-insulating state is destroyed and the character of the fermionic state changes to a state with finite density imbalance in the steady state. Taking fluctuations in the coupling perturbatively into account we still find a bistability, i.e. a region of pumping strengths for which two or more steady states with a distinct finite photon number exist. However, the nature of this bistability at half filling is fundamentally different from the bistability we introduced in Sec. V.1.4. In the case of half filling the nature of the ground state of the effective model, changes its character due to a closing of the gap to the lowest excitations from a Mott-insulator to a charge density wave. Such bistabilities are typically associated with a dissipative first-order transition Larson et al. (2008b, a); Minganti et al. (2018), therefore, they also present at the T=0T\!=\!0 MF level. However, to our knowledge this is the first time the bistability occurring at the self-organization transition of the half-filled Fermi-Hubbard model coupled to the cavity has been described. Furthermore, we show that by varying the cavity parameters, the detuning $\delta$ and the dissipation strength $\varGamma$, we obtain important deviations between the self-consistent thermal state determined by the fluctuations and the T=0T\!=\!0 MF approach.
@@ V.2.1 Self-ordering transition and bistable solutions
Figure 19: (a) Sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$, (b) inverse temperature $\beta J$, (c) scaled photon number and (d) average number of double occupancies as a function of atoms-cavity coupling $\hbar g/J$ at $\hbar\delta/J\!=\!4$; of a finite size system L=6L\!=\!6 at half filling for stable (blue) and unstable (red) results. The T=0T\!=\!0 MF solution is shown as black solid curve. The parameters used are U/J=2U/J\!=\!2, $\hbar\varGamma/J\!=\!1$. Vertical lines denote the critical coupling $\hbar g_{\text{cr}}/J$ and resonances in the atomic limit. The orange dashed lines are the approximate scalings for U/J=0U/J\!=\!0 [Eqs. (39)].
@@
Refer to caption(d)
Figure 20: (a) Sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (b) inverse temperature $\beta J$, (c) scaled photon number and (d) average number of double occupancies as a function of atoms-cavity coupling $\hbar g/J$ at $\hbar\delta/J\!=\!4$; of a finite size system L=6L\!=\!6 for stable (blue) and unstable (red) results; in the weak tunneling perturbation in the thermodynamic limit [Eqs. (29)] stable (dark green) at half filling. The T=0T\!=\!0 MF solution is shown as a black (stable) or gray (unstable) solid curve. The parameters used are U/J=8U/J\!=\!8 and $\hbar\varGamma/J\!=\!1$. Vertical lines denote the critical coupling $\hbar g_{\text{cr}}/J$ (green dashed). The orange dashed line is the approximate scaling for U/J=0U/J\!=\!0 [Eqs. (39)].
To start we discuss the results of the T=0T\!=\!0 MF results. We show an example of the self-organization transition for intermediate interaction strength U/J=2U/J\!=\!2 in Fig. 19 and for larger interaction strength U/J=8U/J\!=\!8 in Fig. 20, where we show with the black line the result of the T=0T\!=\!0 MF method. Within the T=0T\!=\!0 MF method, the photon number, the density imbalance and the average value of the double occupancies exhibit a large value in the self-organized solution at the transition for both shown values of $U/J$. This stems from the property that at commensurate filling a finite density imbalance coincides with the creation of double-occupancies on the low-potential sublattice. At larger interaction strength, U/J=8U/J\!=\!8, the emergence of a bistable behavior around the self-ordering transition can be seen in Fig. 20 for T=0T\!=\!0 MF. The stable solution with homogeneous density distribution and zero photon field and the self-ordered steady state above a critical coupling $g^{\text{MF}}_{\text{cr}}$ (black points) coexist and are connected by an unstable branch (gray points). For the plot shown here, the gray line is partly covered by the red curve. The sublattice imbalance and the double occupancy approach their maximal value and the photon number continues to rise very steeply with the pump strength. At intermediate interaction strength U/J=2U/J\!=\!2 a bistability also appears to occur, but the region of the instability seems to be much more narrow (Fig. 19). Due this very narrow bistable region it becomes numerically difficult to identify it unambiguously.
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
@@
Figure 21: Simplified phase diagrams varying the atoms-cavity coupling $\hbar g/J$ and (a) the on-site interaction $U/J$ at $\hbar\delta/J=4$, $\hbar\varGamma/J\!=\!1$; (b) the pump-cavity detuning $\hbar\delta/J$ at U/J=2U/J\!=\!2, $\hbar\varGamma/J\!=\!1$; (c) the dissipation rate $\hbar\varGamma/J$ at $\hbar\delta/J\!=\!4$, U/J=8U/J\!=\!8 of a finite size system L=6L\!=\!6 at half filling. The lower black line marks the transition between normal [purple region] and the high-$\lambda$ solution in the self-organized phase [blue region] determined by solving Eq. (5) and Eq. (15) with ED, $\hbar g_{\text{cr}}$. At half filling it coincides with the onset of the bistability $\hbar g_{\text{bi,1}}$. The black hatched region marks the region with two stable solutions between $\hbar g_{\text{bi,1}}$ and $\hbar g_{\text{bi,2}}$. The lower red line marks the critical coupling $\hbar g^{\text{MF}}_{\text{cr}}$ obtained for the T=0T\!=\!0 MF approach that coincides with $\hbar g^{\text{MF}}_{\text{bi,1}}$. The red hatched region marks the bistability between $\hbar g^{\text{MF}}_{\text{bi,1}}$ and the upper red line $\hbar g^{\text{MF}}_{\text{bi,2}}$. The dashed lines are approximations of the critical couplings $g_{\text{cr}}^{\text{MF}}$ [Eq. (42)] (magenta dashed) and $\hbar g^{\text{MF}}_{\text{bi,2}}$ (green dashed), determined by solving Eq. (5) at maximal imbalance.
@@
The narrowing at small interaction strength of the parameter region where two stable solutions are present can be seen varying the on-site interaction $U/J$ and atoms-cavity coupling $\hbar g/J$ in Fig. 21 (a), marked for the T=0T\!=\!0 MF results by the red hatched region. At small interaction strength it is very hard to identify whether a bistability is present. In contrast, at large interaction strength a broad bistable region is present. We would like to emphasize that the nature of the bistabilty at half filling differs from the fluctuation-induced bistability at low filling. The onset of the bistability here coincides with the self-organization transition. We attribute this to the change of the nature of the ground state of the effective ionic Hubbard model from the Mott-insulating state to a charge density wave Larson et al. (2008a). This transition we expect around $2\hbar g\lambda\!\approx\!U$ for large $U$ and large $g\lambda$ (as discussed in Sec. IV), where $\lambda$ corresponds to the stable self-ordered solution. The self-consistency condition assuming maximal density imbalance is given by $\langle\hat{\Delta}\rangle\!=\!N$. The resulting critical coupling scales approximately as
@@
(magenta dashed line in Fig. 21). Here, the dependence on the detuning and the dissipation enters as already seen at lower fillings. The approximate scaling with the $\sqrt{U}$ agrees roughly with the scaling obtained numerically in Fig. 21 (a). For the commensurate filling we cannot make the assumption of a small cavity field. Therefore, the considerations made for the critical coupling strength at lower fillings in Sec. V.1.1 do not apply here.
Another approach to determine the bistability edges involves the stability condition derived in Eq. (25). Assuming the solution connecting low-and high field stable solution to be unstable, the coupling for which the condition has a zero with multiplicity 2 is roughly at the two critical couplings. We explored this approach numerically and find reasonable agreement with our previously obtained results from solving Eq. (5) and Eq. (15) via ED.
@@
Figure 22: (a) Sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$, (b) inverse temperature $\beta J$, (c) scaled photon number and (d) average number of double occupancies as a function of $\hbar\delta/J$ at $\hbar g/J\!=\!4$ of a finite size system L=6L\!=\!6 at half filling for stable (blue) and unstable (red) results. The T=0T\!=\!0 MF solution is shown as black solid curve. The parameters used are U/J=2U/J\!=\!2, $\hbar\varGamma/J\!=\!1$. Vertical lines denote the critical coupling $\hbar g_{\text{cr}}/J$ and resonances. The orange dashed lines are the scalings of the analytical approximations for U/J=0U/J\!=\!0 [Eqs. (39)].
@@
Figure 23: (a) Sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (b) inverse temperature $\beta J$, (c) scaled photon number and (d) average number of double occupancies as a function of $\hbar\varGamma/J$ of a finite size system L=6L\!=\!6 at half filling for stable (blue) and unstable (red) results. The T=0T\!=\!0 MF solution is shown as a black (stable) or gray (unstable) solid curve. The parameters used are U/J=8U/J\!=\!8, $\hbar g/J\!=\!6$, $\hbar\delta/J\!=\!4$. The vertical lines denote $2\hbar g\lambda\!=\!U\!+\!\hbar\delta$ (purple dashed) and $\hbar g_{\text{cr}}/J$ (green dashed). The orange dashed line is the scaling of the analytical approximation for U/J=0U/J\!=\!0 [Eqs. (39)].
Refer to caption(a)
Refer to caption(b)
Figure 24: (a) Sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (b) inverse temperature $\beta J$ as a function of atoms-cavity coupling $\hbar g/J$ and on-site interaction $U/J$ of a finite size system L=6L\!=\!6 at half filling. The parameters used are $\hbar\delta/J\!=\!4$, $\hbar\varGamma/J\!=\!1$. The hatched region marks the fluctuation-induced bistability region at the transition to the self-organized phase. The lines denote $U\!=\!\hbar\delta$ (red dashed) and $\hbar g_{\text{cr}}/J$ (green dashed).
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Refer to caption(d)
Figure 25: (a) Sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$, (b) inverse temperature $\beta J$, (c) scaled photon number and (d) average number of double occupancies as a function of atoms-cavity coupling $\hbar g/J$ and detuning $\hbar\delta/J$ of a finite size system L=6L\!=\!6 at half filling. The parameters used are U/J=2U/J\!=\!2, $\hbar\varGamma/J\!=\!1$. The narrow hatched region marks the fluctuation-induced bistability region at the transition to the self-organized phase. The lines denote $U\!=\!\hbar\delta$ (red dashed) and $\hbar g_{\text{cr}}/J$ (green dashed).
@@
We turn to to the dependence with the dissipation strength $\hbar\varGamma/J$. For low values of $\hbar\varGamma/J$, the shift of the self-ordering transition between the T=0T\!=\!0 MF method and the approach including fluctuations is relatively small [black and red hatched region in Fig. 21 (c)]. This is due to the fact that in this regime the self-consistently determined temperature is small compared to the excitation gap, which is proportional to $U/J$. Thus, the finite self-consistent temperature does not introduce a significant shift. However, increasing the dissipation $\hbar\varGamma/J$ the width of the bistability region drastically decreases when increasing $\hbar\varGamma/J$ including the fluctuations, while for the T=0T\!=\!0 MF result it even increases in width [as seen in Fig. 21 (c)]. We attribute the shrinking of the bistability region mostly to the rise in the effective temperature when increasing $\varGamma$ [see blue points in Fig. 23 (b)]. In the presence of the fluctuations, from $\hbar\varGamma_{\text{cr}}/J\!\sim\!18.3$ within the numerical accuracy no bistability can be detected anymore, hinting towards a continuous transition. This result resembles the finding in the model considered in Foss-Feig et al. (2017), where, the ordering transition shows the properties of a discontinuous first-order phase transition which turns continuous in the large dissipation limit.
@@
We comment on how some of the cooling resonances, discussed in detail in Sec. V.1.3, influence the results at half filling. The effects are mostly analogous, i.e. varying $\delta$, we note a drastic increase of the photon number around the cooling resonance $U\!=\!\hbar\delta$ [see Fig. 25 and Fig. 22 (c) around the red dashed line]. Here the energy change when creating a double occupancy is resonant to the process of adding a photon to the cavity field.